OpenStudy (anonymous):

so confused. help? fan and medal (: The equation of line WX is y = −2x − 5. Write an equation of a line perpendicular to line WX in slope-intercept form that contains point (−1, −2). y = 1 over 2x + 3 over 2 y = negative 1 over 2x + 3 over 2 y = 1 over 2x − 3 over 2 y = − 1 over 2x − 3 over 2

OpenStudy (anonymous):

y = 1/2 x + c -2 = -1/2 + c c = -1/2 Equation is y = 1/2 x - 1/2

OpenStudy (anonymous):

If a line is perpendicular to the other one whose slope k, in this case, is -2, the slope of the line is the negative inverse of k, -1/k. So the line function is d=1/2*x+c; Since it passes (−1, −2), substitute it into the function, -2=-1/2+c. c=-3/2. So the perpendicular line is d=1/2*x-3/2